U Brew blends a Colombian coffee worth $0.93 per pound with Kenyan coffee worth $1.20 per pound to get 30 pounds of a coffee blend worth $1.02 per pound. How many pounds of each type of coffee are used
Answer by richwmiller(17219)

x+y=30,
.93x+1.2y=30*1.02
x=20., y=10.